OpenStudy (anonymous):

Which of the following shows the equation log4 64 = 3 written in exponential form? 3^4 = 64 3^64 = 4 4^3 = 64 4^64 = 3

OpenStudy (anonymous):

\(\log_ab=c\iff a^c=b\)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

the base of the log is the base of the exponent the thing it's being set equal to is the power the base is being raised to the thing after the log is what the base raised to the exponent is

OpenStudy (johnweldon1993):

Apply what is shown to your expression \[\large log_a b = c \rightarrow a^c = b\] Compared to your \[\large log_4 64 = 3 \rightarrow ??\]

OpenStudy (anonymous):

is it 4^3=64

OpenStudy (anonymous):

yeah

OpenStudy (johnweldon1993):

Good job :)